OpenStudy (anonymous):

Solve -11(8 – x) = x – 10(5 – x). Please show all work

OpenStudy (l):

-11(8-x) = x-10(5-x) -88+11x = x - 50+10x -88+50 = x+10x-11x -38 = 0 (?) So there is no solution: The solution set is empty set. Because -38 is not equal to zero. There is no x value to satisfy this equation. *